### v3

# FinCoT: Grounding Chain-of-Thought in Expert Financial Reasoning

Abstract
[…]
FinCoT: Grounding Chain-of-Thought in Expert Financial Reasoning
Natapong Nitarach, Warit Sirichotedumrong, Panop Pitchayarthorn, Pittawat Taveekitworachai, Potsawee Manakul, Kunat Pipatanakul SCB 10X, SCBX Group
1 Introduction
Financial decision–making, such as stochastic modeling, risk assessment, portfolio optimization, and algorithmic trading (Markowitz, 1952; Black and Scholes, 1973a; Rockafellar and Uryasev, 2000; Avellaneda and Stoikov, 2008), demands precise mathematics and domain-specific reasoning Lewkowycz et al. (2022); Wen and Zhang (2025). Recent advances in large foundation models for finance, such as the multimodal FinTral (Bhatia et al., 2024) and language-centric Fin-R1 (Liu et al., 2025), demonstrate progress but still face challenges in interpretability and domain alignment (Nie et al., 2024; Arya.ai, 2024; Lee et al., 2025). Accordingly, these shortcomings motivate stricter control over a model’s intermediate reasoning path, which we explore via prompt design.
Prompting guides LLM reasoning without extra training. Methods such as Chain-of-Thought (Wei et al., 2023), Code Prompting (Hu et al., 2023), Plan-and-Solve (Wang et al., 2023), and Self-Reflection (Renze and Guven, 2024) encourage stepwise thinking but remain domain-agnostic. In finance, this can lead to omissions in valuation checks or confusion between basis points and percentages. Yet real-world analysis follows well-defined workflows—valuation, discounting, portfolio attribution—that depend on explicit intermediate structure. Embedding such structure in the prompt helps the model verify units, formulas, and boundary conditions, improving interpretability and alignment with expert practice.
We design FinCoT, a zero-shot prompt that injects expert financial workflows-encoded as Mermaid blueprints-into a structured CoT template, yielding auditable reasoning without fine-tuning. Across ten CFA domains, FinCoT significantly boosts accuracy (most in quantitative areas) and produces shorter, clearer outputs than standard or structured CoT prompts. This paper offers three main contributions:
  • We provide a comprehensive investigation and the first taxonomy of financial prompting–covering standard prompting, unstructured CoT, and structured CoT/FinCoT–clarifying how each paradigm addresses domain-specific reasoning requirements.
  • We release nine blueprint templates—conceptual diagrams modeled after the Unified Modeling Language (UML) (Engels et al., 2000) and rendered in Mermaid syntax–that LLMs can parse as plain–text hints to drive structured reasoning both within FinCoT and in broader domain-specific prompting scenarios.
[…]
Figure 1: Taxonomy of prompting strategies by reasoning structure: SP, UST-CoT, ST-CoT, and FinCoT. Green blocks indicate added reasoning control–CoT steps, tagged thoughts, and expert diagrams (Diagram Embed)–showing the evolution toward more interpretable, domain-aligned prompts.
2 Background and Related Work
2.1 Prompt Engineering
Standard Prompting (SP):
Refers to the baseline technique of simply providing a natural language instruction to an LLM, without providing any intermediate ‘thinking’ steps, demonstrations, or explicit reasoning cues-i.e., a zero-shot setup. While the GPT-3 paper (Brown et al., 2020) popularized few-shot prompting via exemplars, more recent work formalizes and benchmarks zero-shot prompting as a distinct paradigm (Wei et al., 2022). Our implementation follows the formulation shown in Appendix Listing 1 (ZS) in Callanan et al. (2024), and is used to represent the standard prompting baseline.
Unstructured Chain-of-Thought (UST-CoT):
A general-purpose reasoning strategy using free-form CoT to establish a baseline for unconstrained prompting. These include:
• Chain-of-Thought (CoT) (Wei et al., 2023): Decompose reasoning into intermediate steps, encouraging the model to deliberately and systematically ‘think’ before finalizing an answer.
• Code Prompting (Hu et al., 2023): Translates problems into executable code, allowing the model to simulate logic or perform precise computations. In other words, LLMs are elicited to reason explicitly and transparently through code.
• Plan-and-Solve (Wang et al., 2023): Separates planning from solving, where the model first outlines a high-level plan, then executes the reasoning based on that plan.
In addition to these, other prompting techniques have emerged, such as Tree of Thoughts (ToT) (Yao et al., 2023), which explores multiple reasoning paths via tree-structured search; Graph of Thoughts (GoT) (Besta et al., 2024), which frames reasoning as a graph with LLM-generated nodes and edges for flexible information flow. These methods enhance expressiveness for general tasks; they are not tailored for finance, which requires mathematical precision and domain-specific constraints. We adopt the template from Appendix Listing 2 CoT Callanan et al. (2024) as the default prompt for this baseline.
Structured Chain-of-Thought (ST-CoT):
ST-CoT augments a prompt with tags such as <thinking> and <output> that break the model’s response into explicit, modular stages, promoting incremental, easily replaceable reasoning blocks. This tag-driven format has already appeared in open-source trials. [^1] https://gist.github.com/davidmezzetti/1fac6bd406857431f1cdc74545bdfba9 Figure˜1 visually contrasts ST-CoT with SP, UST-CoT, and FinCoT.
FinCoT (§3) inherits ST-CoT’s structure but injects domain-specific Mermaid blueprints to ground each step in expert workflows. Unlike Universal Self-Adaptive Prompting (Wan et al., 2023), which derives few-shot exemplars from LLM memory, FinCoT encodes human-crafted financial reasoning, favoring interpretability and control. The three categories Standard Prompting (direct instruction), Unstructured CoT (free-form reasoning), and Structured CoT (tag-driven with optional expert hints) offer a unified lens for classifying financial prompting.
2.2 LLMs in Domain-Specific Financial Reasoning
Existing approaches to adapting large foundation models for financial reasoning currently fall into three broad paradigms:
Prompting-based:
Methods use few-shot prompts with CFA-style queries, as in “Can GPT Models be Financial Analysts?” Callanan et al. (2024), which evaluate ChatGPT and GPT-4 on mock CFA exams.
Fine-tuning-based:
Methods adapt models via supervised fine-tuning with curated QA/classification datasets (Ma et al., 2023; Harsha et al., 2025) or continued pretraining on domain-specific corpora (Yang et al., 2020; Lee et al., 2024; Bhatia et al., 2024; Ke et al., 2025; Liu et al., 2025). While effective, these approaches rely on labeled data and general language modeling objectives, lacking structuring of financial reasoning, thus limiting interpretability and alignment with expert logic.
Despite advances, prior work has largely overlooked structured reasoning grounded in real-world workflows. We introduce a domain-based prompting framework designed to reflect step-by-step expert logic and evaluate it on CFA-style exam tasks.
3 FinCoT: Augmenting CoT with Structured Financial Expertise
We introduce FinCoT (Financial Chain-of-Thought), a structured prompting framework that enhances LLM reasoning in specialized financial domains. Building upon ST-CoT approaches, FinCoT explicitly embeds expert-derived problem-solving methodologies directly into prompts, guiding LLMs to follow domain-specific reasoning pathways without requiring model fine-tuning. Figure˜2 illustrates the FinCoT architecture, which integrates expert-guided reasoning layers and reflective validation to improve performance in financial tasks.
Figure 2: Overview of FinCoT, integrates structured, expert-guided reasoning layers Mermaid diagrams, plan-and-solve scaffolding, and reflective validation to improve performance in financial tasks.
3.1 The FinCoT Prompt Framework
The FinCoT prompting framework integrates the following key components and logical steps:
  1. 1.
    System: A single, top-level message that frames the task (e.g., “You are a CFA candidate; treat the following as a finance question”).
  2. 2.
    Guided Step-by-Step Execution: The prompt reserves two tag blocks <thinking> for intermediate reasoning and <output> for the final answer-thereby enforcing a structured chain-of-thought (ST-CoT) in one turn.
  3. 3.
    Expert Reasoning Blueprint (via Mermaid Diagram (Sveidqvist and contributors, 2025): A domain-specific, embedded expert blueprint with Mermaid diagram concerning the generation of diagrams (see §4), serve as a "Hint" within the context of the prompt. This blueprint explicitly outlines the recommended step-by-step problem-solving strategy for the given financial domain and is curated through a systematic process detailed in section˜3.2 to ensure consistency and domain alignment.
  4. 4.
    Re-Reading & Semi Self-Reflection: Inside the <output> tags, the model briefly checks its reasoning for consistency before committing the final answer. We call this “semi-reflection” because we drop the separate <reflection> block-avoiding per-step scoring and self-bias noted by Xu et al. (2024) yet still include a short self-check within <output>.
3.2 Embedding Expert Blueprints
The creation of effective expert reasoning blueprints involves a systematic multistage process designed to transform a wide range of financial knowledge into structured and actionable LLM diagrams.
Figure 3: Pipeline for curating financial expert reasoning. Each stage systematically transforms raw financial knowledge into structured Mermaid diagrams for FinCoT prompting.
Curation Pipeline for Expert Reasoning:
To construct expert blueprints for each financial domain, we implement a hybrid pipeline combining machine assistance and human judgment, as illustrated in Figure˜3. The process includes the following stages:
  1. 1.
    Scope Definition and Knowledge Aggregation: The target CFA domain is scoped (e.g., Economics focusing on supply and demand), and relevant expert strategies are aggregated, using Deep Research [^2] An AI agent for retrieving/synthesizing knowledge from public sources such as CFA notes, academic texts, and industry guides., from diverse authoritative sources. Outputs are validated by human-in-the-loop reviewers with financial knowledge (e.g., finance graduates or CFA charterholders) to ensure conceptual accuracy and domain alignment.
  2. 2.
    Validation and Synthesis: We cross-reference and synthesize the aggregated knowledge to ensure accuracy, identify core principles and filter redundancies.
  3. 3.
    Iterative Refinement into Structured Workflows: The synthesized expert knowledge is iteratively transformed into logical step-by-step reasoning workflows. This refinement process focuses on ensuring the coherence, correctness, and clarity of the resulting problem-solving strategies for each financial domain.
  4. 4.
    Mermaid Diagram Generation: This refined workflow is translated into a Mermaid diagram (Bari et al., 2024) using its text-based syntax. We selected Mermaid due to its LLM prompt compatibility and clear visual guidance aligning with the FinCoT prompt. The diagrams are constructed based on the source content validated and synthesized first in 2, and then applied to each financial domain, with the entire collection organized and described in Appendix˜A as reference blueprints [^3] The resulting expert blueprints are reviewed for conceptual consistency and practical correctness (but not guaranteed precision) by CFA Level III charterholders..
  5. 5.
    Prompt Integration: The text-based Mermaid blueprint is embedded as “Hint" within the FinCoT prompt template (Appendix B.2), directly guiding the LLM’s reasoning.
4 Experimental Setup
Model Configurations and Inference Parameters:
We selected the Qwen family of models due to their strong baseline performance in zero-shot financial reasoning. In preliminary evaluations, Qwen2.5-7B-Instruct achieved 69.7% accuracy on standard task prompts, substantially outperforming Llama3.1-8B-Instruct (46.3%), motivating its use as our primary model family. To evaluate the impact of both instruction tuning and domain-specific adaptation, we compare two model groups. (A) General-purpose foundation models
  • Qwen2.5-7B (pretrained model) vs. Qwen2.5-7B-Instruct: to assess the effect of instruction tuning on a strong general-purpose foundation model.
  • Qwen3-8B-Base (pretrained model) vs. Qwen3-8B, Qwen3-8B (Thinker): to isolate the impact of ST-CoT prompting within the same architecture.
  • Gemma-3-12B-IT (Text-only): an instruction-tuned model recently released, comparable in size to Qwen3-8B. It achieved 52.81% accuracy on the Flare-CFA benchmark, outperforming Llama3.1-8B-Instruct, and serves as a competitive baseline.
(B) Financial-specific reasoning models
  • Fin-R1 (7B): adapted from Qwen2.5-7B-Instruct using supervised and reinforcement learning on a financial dataset distilled from DeepSeek-R1 (Liu et al., 2025).
  • DianJin-R1-7B: fine-tuned from Qwen2.5-7B-Instruct using CFLUE, FinQA, and CCC, with GRPO to improve domain-specific reasoning (Zhu et al., 2025).
  • Fin-o1-8B: built on Qwen3-8B and trained on the FinCoT [^4] The FinCoT dataset is constructed by TheFinAI and publicly available at https://huggingface.co/datasets/TheFinAI/FinCoT. It combines financial QA datasets such as FinQA, ConvFinQA, TATQA, DocMath-Eval, Econ-Logic, BizBench-QA, and DocFinQA, with GPT-4o-generated reasoning traces to enhance structured financial question answering. Note that this dataset is not derived from our prompting approach. corpus using SFT and GRPO, setting a strong benchmark in financial reasoning (Qian et al., 2025).
All experiments used a maximum sequence length of 16.384k tokens. Following best practices for decoding stability (Du et al., 2025), we set the generation temperature to 0.2 to encourage focused and consistent outputs under evaluation conditions.
Prompting Strategies Compared:
Our study evaluates the effectiveness of FinCoT against three baseline prompting strategies: SP, UST-CoT, and ST-CoT, which were detailed in Section˜2. For clarity in this section:
  • SP: The model receives only the target question.
  • UST-CoT: In addition to the question, the model is given a generic cue to reason step-by-step.
  • ST-CoT: This strategy employs structural tags (e.g., <thinking>, <output>) to guide the model in generating an organized step-by-step reasoning trace for the target question.
  • FinCoT: A zero-shot prompting method that integrates expert domain templates. Each prompt includes a Mermaid diagram as a “Hint” to guide structured financial reasoning via relevant domain insights.
While all methods operate in a zero-shot setting, FinCoT uniquely injects expert-guided structure through diagrams. Recent studies suggest that even large reasoning models may struggle with instruction-following when overloaded with reasoning cues (Li et al., 2025; Fu et al., 2025; Jang et al., 2025; Yao et al., 2025), though this remains underexplored in financial contexts. We thus evaluate whether CoT-style prompts (UST-CoT, ST-CoT, FinCoT) enhance instruction-following compared to SP.
Evaluation Benchmark:
To assess financial reasoning, we use 1.032k multiple-choice questions from the CFA-Easy subset of FinEval (also referred to as Flare-CFA), originally introduced by Ke et al. (2025). This curated set reflects the rigor of CFA exams and enables evaluation across both theoretical and practical domains. Each question is categorized into one of ten CFA domains using GPT-4o with a dedicated classification prompt (see Appendix B.3), and Figure C shows the resulting domain distribution.
Evaluation Metrics:
We report accuracy as the metric, defined as the percentage of questions where the model’s prediction matches the ground truth. To assess response efficiency, we also report the average output length (in tokens) across questions. For statistical significance, we use a paired bootstrap test (Efron and Tibshirani, 1994) with B=10kB{=}10k resamples over binary correctness scores, reporting the mean difference ($\Delta$), 95% confidence interval, and $p$-value. Additionally, we measure the proportion of financial domains where a method improves accuracy by at least 1% over SP and compute the average domain-wise accuracy gain.
5 Results and Discussions
5.1 Financial Reasoning Performance
Prompt Accuracy (%)
General models Financial models
Qwen2.5-7B
Qwen2.5-7B
Instruct
Qwen3-8B
Base
Qwen3-8B
Qwen3-8B
Gemma-3-12B
IT
Fin-R1
7B
DianJin-R1
7B
Fin-o1
8B
SP 54.07 69.67 63.18 74.42 88.18 52.81 65.70 78.39 79.65
UST-CoT 67.83 ($\uparrow$13.76) 75.68* ($\uparrow$6.01) 72.58 ($\uparrow$9.40) 82.36* ($\uparrow$7.94) 89.05* ($\uparrow$0.87) 77.81* ($\uparrow$25.00) 75.19 ($\uparrow$9.49) 67.73 ($\downarrow$10.66) 79.36 ($\downarrow$0.29)
ST-CoT 70.35* ($\uparrow$16.28) 74.52 ($\uparrow$4.85) 78.49 ($\uparrow$15.31) 81.01 ($\uparrow$6.59) 88.18 76.74 ($\uparrow$23.93) 74.32 ($\uparrow$8.62) 68.80 ($\downarrow$9.59) 78.39 ($\downarrow$1.26)
FinCoT 62.02 ($\uparrow$7.95) 74.22 ($\uparrow$4.55) 80.52* ($\uparrow$17.34) 81.10 ($\uparrow$6.68) 87.21 ($\downarrow$0.97) 75.58 ($\uparrow$22.77) 75.78* ($\uparrow$10.08) 79.75* ($\uparrow$1.36) 77.23 ($\downarrow$2.42)
Domain-wise performance of FinCoT
Economics 69.09 ($\uparrow$15.02) 73.26 ($\uparrow$3.59) 79.26 ($\uparrow$16.08) 79.55 ($\uparrow$5.13) 86.92 ($\downarrow$1.26) 74.61 ($\uparrow$21.80) 73.45 ($\uparrow$7.75) 55.52 ($\downarrow$22.87) 78.00 ($\downarrow$1.65)
FixedIncome 68.12 ($\uparrow$14.05) 73.35 ($\uparrow$3.68) 78.88 ($\uparrow$15.70) 80.81 ($\uparrow$6.39) 87.21 ($\downarrow$0.97) 76.45 ($\uparrow$23.64) 74.22 ($\uparrow$8.52) 66.86 ($\downarrow$11.53) 76.74 ($\downarrow$2.91)
Quant.Meth. 68.02 ($\uparrow$13.95) 75.19 ($\uparrow$5.52) 80.14 ($\uparrow$16.96) 80.91 ($\uparrow$6.49) 87.79 ($\downarrow$0.39) 75.68 ($\uparrow$22.87) 74.90 ($\uparrow$9.20) 65.79 ($\downarrow$12.60) 77.42 ($\downarrow$2.23)
EquityInvest. 69.09 ($\uparrow$15.02) 74.22 ($\uparrow$4.55) 79.26 ($\uparrow$16.08) 80.52 ($\uparrow$6.10) 86.72 ($\downarrow$1.46) 76.45 ($\uparrow$23.64) 74.42 ($\uparrow$8.72) 62.31 ($\downarrow$16.08) 78.68 ($\downarrow$0.97)
Port.Manage. 67.54 ($\uparrow$13.47) 74.13 ($\uparrow$4.46) 80.72 ($\uparrow$17.54) 80.91 ($\uparrow$6.49) 86.92 ($\downarrow$1.26) 77.03 ($\uparrow$24.22) 75.00 ($\uparrow$9.30) 62.02 ($\downarrow$16.37) 76.55 ($\downarrow$3.10)
Derivatives 68.90 ($\uparrow$14.83) 73.64 ($\uparrow$3.97) 79.84 ($\uparrow$16.66) 80.81 ($\uparrow$6.39) 87.21 ($\downarrow$0.97) 77.23 ($\uparrow$24.42) 76.16 ($\uparrow$10.46) 71.80 ($\downarrow$6.59) 79.94 ($\uparrow$0.29)
Fin. Reporting 69.28 ($\uparrow$15.21) 73.84 ($\uparrow$4.17) 79.07 ($\uparrow$15.89) 81.10 ($\uparrow$6.68) 87.02 ($\downarrow$1.16) 75.87 ($\uparrow$23.06) 72.87 ($\uparrow$7.17) 62.69 ($\downarrow$15.70) 77.23 ($\downarrow$2.42)
Alter.Invest. 68.99 ($\uparrow$14.92) 74.90 ($\uparrow$5.23) 78.97 ($\uparrow$15.79) 79.94 ($\uparrow$5.52) 87.50 ($\downarrow$0.68) 76.94 ($\uparrow$24.13) 74.90 ($\uparrow$9.20) 56.98 ($\downarrow$21.41) 78.88 ($\downarrow$0.77)
Corp.Issuers 68.31 ($\uparrow$14.24) 74.32 ($\uparrow$4.65) 79.26 ($\uparrow$16.08) 79.36 ($\uparrow$4.94) 87.02 ($\downarrow$1.16) 77.23 ($\uparrow$24.42) 75.58 ($\uparrow$9.88) 60.08 ($\downarrow$18.31) 79.07 ($\downarrow$0.58)
Table 1: Comparison of accuracy (%) of prompting techniques. ‘FinCoT’ simultaneously applies expert reasoning blueprints from all CFA domains, while each ‘(DomainName)’ (e.g., ‘Economics’) row applies domain-specific blueprints individually. ($\uparrow$$\downarrow$) Denote accuracy improvement or decline relative to the SP baseline, colored green for ($\uparrow$) and red for ($\downarrow$). Bold values highlight the best-performing prompt variant for each model. (*) Indicates that the accuracy improvement among the model-level prompt variants is statistically significant ($p<0.05$) based on paired bootstrap testing; domain-specific rows are not tested for significance.
Baseline Performance:
Table˜1 reports zero-shot accuracies for four prompting strategies (SP, UST-CoT, ST-CoT, FinCoT) across our model suite. Under the basic SP prompt, the instruction-tuned Qwen3-8B (Thinker) attains the highest accuracy among general-purpose models (88.18%), while the financial model Fin-o1-8B leads its group at 79.65%. These strong baselines clearly highlight the effectiveness of instruction tuning and domain specificity.
Pretrained Models:
On Qwen2.5-7B, FinCoT (All Blueprints) yields a +7.95 pp improvement over SP (95% CI [6.30, 9.59], $p<0.001$), while UST-CoT and ST-CoT perform worse than the baseline. When FinCoT is applied using a single-domain blueprint (e.g., Financial Reporting), the gains increase substantially to +15.21 pp. Similarly, on Qwen3-8B-Base, FinCoT delivers the strongest overall boost (+17.35 pp, 95% CI [15.02, 19.77], $p<0.001$). These findings further underscore the importance of structured domain knowledge, particularly for models lacking instruction or domain alignment.
Instruction Models:
We evaluate three prominent instruction-tuned variants. On Qwen2.5-7B-Instruct, FinCoT improves accuracy by +4.55 pp over SP, compared to -1.46 pp with UST-CoT and -0.3 pp with ST-CoT. On Qwen3-8B (Thinker), FinCoT yields a slight drop (–0.96 pp), while ST-CoT shows no change and UST-CoT yields a modest +0.87 pp. Gemma-3-12B-IT, a strong instruction-tuned baseline (52.81% SP), benefits substantially from all strategies: +25.00 pp (UST-CoT), +23.93 pp (ST-CoT), and +22.77 pp (FinCoT). Notably, domain-specific FinCoT prompts (e.g., Derivatives, Corporate Issuers) provide even larger boosts (+24.42 pp), indicating that blueprint reasoning complements instruction tuning by addressing specialized financial gaps.
Financial-Specific Models:
FinCoT also helps specialized models like Fin-R1, confirming that blueprint prompting provides complementary gains beyond fine-tuning. However, for models with strong built-in reasoning, such as DianJin-R1-7B and Fin-o1-8B, FinCoT offers limited improvement or slight degradation—likely due to conflicts between external scaffolds and internal reasoning routines. These outcomes suggest diminishing returns for CoT prompting when domain alignment and reasoning are already deeply encoded.
Overall, FinCoT is most impactful for models lacking prior task-specific adaptation. By grounding reasoning in structured financial workflows, it bridges key gaps in zero-shot settings without requiring additional tuning. This pattern highlights a trade-off between model internalization and prompt-time controllability. Future work could explore hybrid strategies that adapt prompting depth based on model alignment.
Figure 4: Accuracy improvements (%) of each FinCoT domain-specific prompt compared to Standard Prompting (SP). Subfigure (a) shows results on Qwen3-8B-Base (pretrained), while (b) shows results on Fin-o1-8B (finance-specific).
Cross-domain behavior of FinCoT:
This section examines how pretrained and finance-specific models respond to structured prompting with FinCoT. Each domain-specific blueprint is applied across all CFA domains to evaluate its transferability, and accuracy differences relative to SP are measured. Figure 4(a) and 4(b) visualize results for Qwen3-8B-Base (pretrained) and Fin-o1-8B (finance-specific), while Table 1 provides a comprehensive summary of overall model-level accuracy. On Qwen3-8B-Base, FinCoT consistently improves performance. Prompts from quantitative domains such as Derivatives, Portfolio Management, and Corporate Issuers yield average gains exceeding +13 pp. The blueprint structure provides inductive guidance that enhances decomposition, formula selection, and financial term alignment. In several domains, Qwen3-8B-Base with FinCoT matches or surpasses the SP baseline of Fin-o1-8B, despite the absence of any task-specific training.
On Fin-o1-8B, gains from FinCoT are more modest, typically within the +1–4 pp range. Minor declines appear in some domains (e.g., Fixed Income, Equity Investments), suggesting that additional scaffolding may interfere with optimized internal reasoning acquired during fine-tuning. Structured prompts may over-specify solutions or reduce instruction-following flexibility.
These findings highlight FinCoT’s complementary role. For pretrained models, FinCoT acts as a lightweight yet effective augmentation layer at inference time, reducing the performance gap with fine-tuned models. For already fine-tuned models, careful prompt selection or adaptation may be necessary to preserve existing reasoning strengths without introducing conflict. A broader breakdown of FinCoT performance across additional models is provided in Appendix F, where radar plots illustrate domain-wise patterns and complement the main analysis.
Figure 5: Comparison of prompt strategies across financial domains. Each point represents a domain-method pair, with position indicating accuracy improvement and domain coverage. Circle size encodes generated tokens.
5.2 Efficiency Analysis
Effective deployment of foundation models in financial settings requires not only accuracy but also efficiency in output length and computational cost. This section investigates the verbosity and accuracy trade-offs across prompting strategies. While Fin-o1-8B, DianJin-R1-7B, and Fin-R1 (7B) are fine-tuned models optimized for financial reasoning, FinCoT offers a prompting-based alternative that achieves competitive performance without additional training. Comparative accuracy is illustrated in Figure 5, and a detailed breakdown of output lengths across models and domains is provided in Appendix D (Table 2) to support efficiency analysis.
Output Token Length vs. Accuracy:
In practical deployments, prompting strategies that achieve high accuracy with minimal output length are especially desirable. This section compares the token efficiency of FinCoT with standard prompting strategies, focusing on general-purpose and financial-specific models.
[…]
Three behavioral zones emerge from these results. First, models like Gemma-3-12B-IT and Qwen3-8B-Base show high mean improvement and strong cross-domain transfer, reflecting effective generalization. Second, models such as Qwen2.5-7B-Instruct and Qwen3-8B display noticeable cross-domain transfer with lower mean improvement, suggesting limited benefit. Third, models like Fin-o1-8B and Qwen3-8B (Thinker) exhibit low cross-domain adaptability and minimal performance lift, with signs that overly detailed prompting may interfere with internal reasoning.
These findings highlight the importance of aligning the prompting strategy with a model’s pretraining or fine-tuning characteristics to optimize both performance and efficiency. With limited supervision, FinCoT offers a transparent, cost-effective alternative for enhancing financial reasoning.
6 Conclusion
We presented FinCoT, a zero-shot prompting framework that embeds expert-curated Mermaid diagrams within structured chain-of-thought scaffolds. By grounding reasoning in domain logic, FinCoT bridges human financial workflows with LLM outputs, without model fine-tuning. While broadly applicable, FinCoT yields strong gains for general models (e.g., Qwen3-8B-Base), but more modest or negative effects for instruction or finance-specific models (e.g., Qwen-Thinker, Fin-o1, DianJin-R1), where added structure may interfere with learned reasoning.
Relative to SP, FinCoT improves Qwen3-8B-Base by +17.33 pp and Fin-R1 by +10.08 pp ($p<0.001$), outperforming some fine-tuned models. In contrast, instruction-tuned models like Qwen3-8B (Thinker) sometimes favor UST-CoT. Cross-domain results show blueprints from quantitative fields transfer best (+27.3 pp on Qwen3-8B-Base). FinCoT also reduces token output by up to $8\times$, offering interpretable, efficient prompting for regulated financial applications. Ultimately, FinCoT suggests that with meticulous prompt design, even general-purpose LLMs can approach the reasoning quality of fine-tuned financial experts in complex decision-making tasks.
Limitations
(i) Blueprint creation currently requires expert review ($\sim$2 hours/domain); (ii) evaluations are limited to multiple-choice settings; (iii) interpretability is measured using rubric scores, not human studies.
Future work includes automating blueprint generation and integrating retrieval-augmented diagrams to expand FinCoT’s applicability.
Our findings suggest that expert-guided prompting acts as a lightweight form of knowledge distillation, embedding structured reasoning in LLM outputs without altering model weights. This method may generalize to high-stakes fields like law, medicine, and engineering, and underscores the value of prompt-level supervision for interpretable, efficient domain reasoning.
Acknowledgments
We extend our gratitude to the Hatari Team for generously providing additional GPU resources, which significantly enhanced our computational capacity and enabled the successful execution of large-scale experiments. We also thank Tuntai Boriboonthana (CFA Charterholder) from InnovestX for his expert contributions in verifying blueprint fidelity and conducting random audits to ensure domain classification consistency in GPT-4o outputs. Their support and insights were invaluable to the development and evaluation of this work.
Appendix A Expert Reasoning Blueprints
Economics
***Economics:***
‘‘‘mermaid
graph TD;
  A[Step 1: Question Breakdown] -->|Extract key terms| A1{Identify Topic}
  A1 -->|Micro: Supply & Demand, Market Structures| A2
  A1 -->|Macro: GDP, Growth, Policy, Trade| A3
  A1 -->|Currency & Regulation| A4
  A2 --> B1[Identify model: Elasticity, Cost Curves, Shutdown Points]
  A3 --> B2[Map to AD-AS, Business Cycles, Growth Theories]
  A4 --> B3[Assess Exchange Rates, Trade, Capital Flows, Regulation]
  B1 -->|Check for formula or concept?| C{Numerical or Conceptual}
  B2 --> C
  B3 --> C
  C -->|Numerical| D1[Extract data, apply formulas, check assumptions]
  C -->|Conceptual| D2[Analyze cause-effect, policy impact]
  D1 --> E[Step 4: Solution Development]
  D2 --> E
  E -->|Construct structured response| E1(Core insight + economic rationale)
  E -->|Consider alternative scenarios| E2(Assess different possibilities)
  E1 --> F[Step 5: Answer Validation]
  E2 --> F
  F -->|Check logic, principles, and assumptions| F1(Verify consistency)
  F1 -->|Ensure completeness & clarity| F2(Confirm answer structure)
Explanation:
Step-1: Question Breakdown (A) – Extract key terms by parsing the question to see whether it focuses on microeconomics, macroeconomics, or currency/regulation topics (300Hours, 2025a; UWorld Finance, 2025a).
Step-2: Identify Topic (A1) – Microeconomics (A2): Focus on supply & demand mechanisms and market structures such as perfect competition, monopoly, oligopoly, and monopolistic competition (Investopedia, 2025a; 300Hours, 2025a). – Macroeconomics (A3): Consider aggregate demand–aggregate supply analysis, phases of the business cycle (expansion, peak, contraction, trough), and growth models (Solow, endogenous growth) (Investopedia, 2025b; CFA Institute, 2024). – Currency & Regulation (A4): Examine exchange-rate regimes (floating vs. pegged), trade balances, capital-flow impacts, and relevant government policies (Investopedia, 2025, 2025b).
Step-3: Model Selection or Strategy Mapping (B1–B3) – Micro Models (B1): Choose elasticity calculations and cost-curve analysis (marginal/average cost, shutdown point) for supply–demand or firm-behavior questions (Investopedia, 2025e). – Macro Frameworks (B2): Apply AD–AS curves, Phillips-curve trade-offs, or business-cycle indicators to frame policy or growth analysis (Investopedia, 2025, 2025c). – FX & Regulation (B3): Use exchange-rate determination models, balance-of-payments analysis, or regulatory impact frameworks for currency/trade questions (Investopedia, 2025a).
Step-4: Determine Numerical vs. Conceptual Approach (C) – Numerical (D1): Gather the relevant data (prices, quantities, rates), apply formulae (e.g., $\text{Elasticity}=\frac{\%\Delta Q_{n}}{\%\Delta P}$, $\text{GDP}=C+I+G+(X-M)$ ), and verify assumptions (Investopedia, 2025c). – Conceptual (D2): Construct a narrative explaining cause–effect relationships (e.g., how a monetary-policy change shifts AD or how trade barriers affect capital flows) (Investopedia, 2025b).
Step-5: Solution Development (E) – Structured Response (E1): State the core economic insight first, then provide the step-by-step rationale linking theory to the question context (Kaplan Schweser, 2025). – Alternative Scenarios (E2): Where relevant, outline best-case, base-case, and worst-case scenarios or show how supply–demand curves shift under different assumptions (iPassFinanceExams, 2025).
Step-6: Answer Validation (F) – Verify Consistency (F1): Check that numerical answers satisfy boundary conditions (e.g., correct sign on elasticity, GDP component sums). – Confirm Clarity (F2): Ensure your explanation is complete, logically ordered, and clearly communicates both the result and its limitations (UWorld Finance, 2025b).
Source:
300Hours CFA Level 1 Economics Cheat Sheet (300Hours, 2025a), UWorld Finance’s CFA® Economics: Syllabus & Sample Questions (UWorld Finance, 2025a), Kaplan Schweser’s Level I Economics tips (Kaplan Schweser, 2025), Efficient Learning’s CFA Economics overview (Efficient Learning, 2025), iPass Finance Exams’ study guide (iPassFinanceExams, 2025), AnalystPrep’s essential review summary (AnalystPrep, 2025), and key Investopedia articles on supply and demand (Investopedia, 2025a), GDP (Investopedia, 2025c), and business cycles (Investopedia, 2025b) as well as Prof. Brian Gordon’s CFA Exam Level 1 Economics video (Gordon, 2020a).
Fixed Income
***Fixed Income:***
‘‘‘mermaid
graph TD
    A[Purpose and Scope] --> B3[Analyze Macro Conditions]
    B --> C[Assess Bond Features]
    C --> D[Risk and Yield Analysis]
    D --> E[Develop Recommendations]
    E --> F[Review Performance]
    %% Notes and detailed steps
    A --> |Set objectives| B
    B --> |Review interest rates and inflation| C
    C --> |Focus on duration, spread| D
    D --> |Assess scenarios| E
Explanation:
Step-1: Purpose and Scope – Define the investment objective–income generation, capital preservation, hedging, or total return and establish portfolio constraints and benchmarks such as target yield, duration limits, credit quality floors, or sector allocation guidelines (Investopedia, 2025b).
Step-2: Analyze Macro Conditions – Examine current and forecast interest rate paths, since rising rates erode bond prices and falling rates support them (Investopedia, 2025); monitor inflation indicators (CPI, PPI) to gauge real yield trends (Investopedia, 2025a); and assess yield-curve shapes (normal, inverted, flat) for economic turning points and yield-curve trade opportunities (Investopedia, 2025b).
Step-3: Assess Bond Features – Identify bond type government, corporate, municipal, structured products (ABS/MBS) and note any embedded options (callable, putable, convertible) (Investopedia, 2025); review coupon structure (fixed vs. floating), payment frequency, and maturity to understand cash-flow timing and reinvestment risk (Investopedia, 2025c).
Step-4. Risk and Yield Analysis – Calculate duration to estimate price sensitivity to yield changes (Investopedia, 2025e) and convexity for non-linear price effects (Investopedia, 2025b); analyze credit spreads over benchmarks to gauge default and liquidity risk (Investopedia, 2025d); and stress-test the portfolio under parallel shifts, steepeners, and flatteners to assess P&L impacts.
Step-5: Develop Recommendations – Formulate strategies such as adjusting overall duration (shorten if rates are likely to rise), implementing barbell or laddered maturity structures, or choosing bullet portfolios to manage reinvestment and rate risk (James, 2025).
Step-6: Review Performance – Track total returns (price changes plus coupon income) against benchmarks like the Bloomberg US Aggregate Bond Index (Bloomberg, 2025); perform attribution analysis to decompose yield carry, curve roll-down, and spread effects; and revisit assumptions and rebalance when market conditions or issuer fundamentals change (Investopedia, 2025a).
Source:
CFA Program Curriculum for Fixed Income (CFA Institute, 2025e), Fabozzi’s Bond Markets, Analysis, and Strategies (Fabozzi, 2012), Tuckman & Serrat’s Fixed Income Securities (Tuckman and Serrat, 2011), Jarrow & Turnbull’s credit-risk derivatives pricing (Jarrow and Turnbull, 1995), Basel Committee papers on credit risk, Investopedia articles on fixed-income concepts (Investopedia, 2025b, b, e), Reuters coverage of convexity risk, and Dichev’s balance sheet model for distress prediction (Dichev, 2017).
Quantitative Methods
***Quantitative Methods:***
‘‘‘mermaid
graph TD
    A["Articulating Purpose and Context"] --> B["Collecting Input Data"]
    B --> C["Processing and Cleaning Data"]
    C --> D["Selecting Quantitative Models and Tools"]
    D --> E["Estimating Parameters and Testing Hypotheses"]
    E --> F["Interpreting Results and Communicating Findings"]
    F --> G["Monitoring and Model Reassessment"]
Explanation:
Step-1: Articulating Purpose and Context (A) Define the research question time value of money calculations, probability distributions, hypothesis testing, regression analysis, or portfolio statistics–and establish the CFA application context (market efficiency, risk estimation, cash-flow forecasting) (CFA Institute, 2025h).
Step-2: Collecting Input Data (B) Gather historical returns, economic indicators, financial statements, and market data from reputable sources; ensure relevance by matching data to the chosen objective (e.g., interest rates for TVM, volatility for risk models) (Investopedia, 2025f).
Step-3: Processing and Cleaning Data (C) Perform data quality checks and remove outliers, handle missing values, confirm consistency–and apply transformations (normalization, log-transforms) before analysis (Wooldridge, 2013).
Step-4: Selecting Quantitative Models and Tools (D) Choose appropriate models–ARIMA for time series, linear/multivariate regression, probability distributions, or Monte Carlo simulation–and leverage CFA-recommended software or spreadsheet tools (Damodaran, 2012; Investopedia, 2025g).
Step-5: Estimating Parameters and Testing Hypotheses (E) Estimate model parameters via regression or maximum likelihood; conduct t-tests, F-tests, or chi-square tests to validate assumptions and results, with Level II emphasis on multivariate regression and sensitivity analysis (Wooldridge, 2013).
Step-6: Interpreting Results and Communicating Findings (F) Translate coefficients, p-values, and confidence intervals into actionable investment insights; prepare clear visual aids (charts, tables) to support recommendations (Bodie et al., 2017).
Step-7: Monitoring and Model Reassessment (G) Track out-of-sample performance against benchmarks; update models as new data arrive, reassess assumptions, and recalibrate parameters to maintain relevance (Wikipedia contributors, 2025).
Source:
CFA Program Curriculum: Quantitative Methods (CFA Institute, 2025h), Wooldridge’s Introductory Econometrics (Wooldridge, 2013), Damodaran’s Investment Valuation (Damodaran, 2012), and Bodie, Kane, & Marcus’s Investments (Bodie et al., 2017).
Equity Investments
***Equity Investing:***
‘‘‘mermaid
graph TD
    A[Objective Setting] --> B[Market and Sector Insights]
    B --> C[Industry Competitive Analysis]
    C --> D[Company Review]
    D --> E[Valuation and Risks]
    E --> F[Investment Decision]
    %% Step-specific highlights
    B --> |Look at growth patterns| C
    C --> |Evaluate competitors’ positions| D
    D --> |Check financial health| E
    E --> |Combine insights into strategy| F
Explanation
Step-1: Objective Setting (A) Define your investment objectives–capital appreciation, dividend income, or total return in line with your risk tolerance and investment horizon; consider constraints such as liquidity needs, tax implications, regulatory requirements, and any specific mandates (CFA Institute, 2025b).
Step-2: Market and Sector Insights (B) Assess macro indicators (GDP growth, interest rates, inflation) to gauge the overall market environment and identify sectors poised for growth or decline based on economic trends, technological shifts, and consumer behavior (Investopedia, 2025a).
Step-3: Industry Competitive Analysis (C) Apply Porter’s Five Forces to evaluate industry attractiveness–competitive rivalry, threat of new entrants, bargaining power of suppliers and buyers, and substitute threats–and assess each firm’s market share and competitive moat (Investopedia, 2025d).
Step-4: Company Review (D) Examine financial statements (income statement, balance sheet, cash flows) to measure profitability, liquidity, and stability; evaluate management’s track record and strategic vision; and review corporate governance structures to ensure alignment with shareholder interests (Investopedia, 2025d; Bodie et al., 2017).
Step-5: Valuation and Risks (E) Use valuation methods–Discounted Cash Flow (DCF), Price-to-Earnings (P/E) ratios, Dividend Discount Models (DDM)–to estimate intrinsic value; identify key risks such as market volatility, operational challenges, regulatory changes, and competitive threats (Pinto et al., 2015).
Step-6: Investment Decision (F) Formulate your Buy, Hold, or Sell recommendation based on the above analyses and determine how the position fits within the broader portfolio–considering diversification, correlation, and overall risk–return objectives (300Hours, 2025b).
Source:
CFA Program Curriculum’s Equity Investments module (CFA Institute, 2025b), Investopedia’s guides on fundamental analysis (Investopedia, 2025a), Porter’s Five Forces stock analysis (Investopedia, 2025d), and reading financial reports (Investopedia, 2025d), 300Hours’ CFA Level 1 Equity Cheat Sheet (300Hours, 2025b), Bodie, Kane & Marcus’s Investments (Bodie et al., 2017), Pinto et al.’s Equity Asset Valuation (Pinto et al., 2015), and CFA Level I Equity video lectures by Prof. Brian Gordon (Gordon, 2020b, c).
Portfolio Management
***Portfolio Management:***
‘‘‘mermaid
graph TD
    A["Define Investment Objectives"] --> B["Establish Investment Constraints"]
    B --> C["Develop Strategic Asset Allocation"]
    C --> D["Incorporate Tactical Adjustments"]
    D --> E["Select and Optimize Securities"]
    E --> F["Execute Implementation and Trading"]
    F --> G["Measure Performance and Attribution"]
    G --> H["Monitor Risk and Compliance"]
    H --> I["Rebalance and Adjust Portfolio"]
Explanation:
Step-1: Define Investment Objectives – Clarify whether the portfolio is aimed at capital growth, income generation, or a balanced mix. Specify expected returns, risk tolerance, and liquidity needs. This step forms the foundation for aligning investment strategy with client mandates (CFA Institute, 2025g).
Step-2: Establish Investment Constraints – Define legal, regulatory, tax, and unique client considerations such as ESG preferences or geographic limits. These constraints ensure feasibility and compliance of portfolio design (CFA Institute, 2025g).
Step-3: Develop Strategic Asset Allocation – Allocate across major asset classes (equities, fixed income, alternatives, cash) based on expected returns and risk tolerance. Use models from Modern Portfolio Theory and CAPM to inform allocation (Markowitz, 1952; Bodie et al., 2017).
Step-4: Incorporate Tactical Adjustments – Introduce short-term adjustments to the strategic allocation based on market outlook or economic indicators. These shifts aim to enhance returns through asset or sector rotation (Grinold and Kahn, 2000).
Step-5: Select and Optimize Securities – Apply quantitative screens and qualitative research to choose securities. Use optimization techniques such as mean-variance optimization or the Black-Litterman model to maximize risk-adjusted returns (Bodie et al., 2017; Grinold and Kahn, 2000).
Step-6: Execute Implementation and Trading – Implement trade strategies that minimize costs and slippage, considering market impact and liquidity. Align execution with strategic intentions (CFA Institute, 2025g).
Step-7: Measure Performance and Attribution – Track performance using return metrics, Sharpe ratio, alpha, and beta. Perform attribution to evaluate decisions across asset allocation, sector, and security selection (Grinold and Kahn, 2000).
Step-8: Monitor Risk and Compliance – Use tools like Value-at-Risk (VaR), stress testing, and tracking error to monitor portfolio risk. Ensure compliance with constraints and regulations (CFA Institute, 2025g).
Step-9: Rebalance and Adjust Portfolio – Periodically adjust the portfolio to maintain alignment with the strategic asset allocation as market conditions evolve.
Source:
CFA Program Curriculum’s Portfolio Management module (CFA Institute, 2025g), Bodie, Kane & Marcus’s Investments for portfolio theory and risk-return optimization (Bodie et al., 2017), Grinold & Kahn’s Active Portfolio Management for advanced attribution and optimization techniques (Grinold and Kahn, 2000), and Markowitz’s seminal Portfolio Selection on diversification and risk-adjusted returns (Markowitz, 1952).
Derivatives
***Derivatives:***
‘‘‘mermaid
graph TD
    A[Define Objective and Context] --> B[Identify Derivative Instrument]
    B --> C[Understand Contract Specifications]
    C --> D[Gather Market Data]
    D --> E[Apply Valuation Models]
    E --> F[Assess Risks: Market, Counterparty, etc.]
    F --> G[Construct Payoff Diagrams or Strategies]
    G --> H[Interpret Results and Make Recommendations]
    H --> I[Review, Monitor, and Adjust Strategies]
    %% Example labels or notes (optional)
    A --> |Hedging, speculation, arbitrage| B
    C --> |Features like notional amount, expiration| D
    D --> |Market prices, volatility, risk-free rates| E
    F --> |Sensitivity to Greeks: Delta, Gamma, Vega, etc.| G
    H --> |Adjust based on changing market conditions| I
Explanation:
Step-1: Define Objective and Context – Clarify the purpose of using derivatives: hedging, speculation, or arbitrage. Identify relevant constraints, such as regulatory limitations or portfolio mandates (CFA Institute, 2025f; Hull, 2017).
Step-2: Identify Derivative Instrument – Choose the appropriate derivative: options, futures, forwards, swaps, or structured/exotic products (Jarrow and Turnbull, 1996).
Step-3: Understand Contract Specifications – Review contract parameters, including the underlying asset, strike price, expiration, settlement method (physical or cash), and style (European, American) (CFA Institute, 2025f).
Step-4: Gather Market Data – Collect input variables such as spot price, volatility, risk-free rate, dividends, and term structure of interest rates (Hull, 2017).
Step-5: Apply Valuation Models – Apply pricing frameworks suited to the derivative:
  • Black-Scholes model for European options (Black and Scholes, 1973b).
  • Binomial Tree for path-dependent or American-style options (Hull, 2017).
  • Cost-of-carry model for futures and forwards (Jarrow and Turnbull, 1996).
  • Finite-difference methods for complex derivatives (Tavella and Randall, 2000).
Step-6: Assess Risks – Use Greeks (Delta, Gamma, Vega, Theta, Rho) to evaluate sensitivity to market factors. Consider counterparty and credit risk in OTC markets (Hull, 2017; Board, 2025).
Step-7: Construct Payoff Diagrams or Strategies – Visualize outcomes using payoff graphs. Design strategies such as straddles, collars, or protective puts based on desired exposure (Hull, 2017).
Step-8: Interpret Results and Make Recommendations – Translate model output into actionable insights: confirm hedge effectiveness, profit potential, or risk exposure.
Step-9: Review, Monitor, and Adjust Strategies – Continuously monitor derivative positions in light of market conditions, risk metrics, and investment objectives (Board, 2025).
Source:
Based on Hull’s comprehensive treatment of markets and pricing models (Hull, 2017), the CFA Institute Level II Derivatives readings (CFA Institute, 2025f), Black & Scholes’s seminal option pricing model (Black and Scholes, 1973b), Jarrow & Turnbull’s practical engineering perspective (Jarrow and Turnbull, 1996), Tavella & Randall’s numerical finite-difference techniques (Tavella and Randall, 2000), and the Basel Committee’s OTC derivatives reforms for regulatory context (Board, 2025).
Financial Reporting
***Financial Reporting:**
‘‘‘mermaid
graph TD
A[Articulating Purpose and Context] --> B[Collecting Input Data]
    B --> C[Processing Data]
    C --> D[Analyzing and Interpreting Processed Data]
    D --> E[Developing and Communicating Conclusions]
    E --> F[Doing Follow-Up]
    A --> |Defines goals, tools, and audience| B
    B --> |Gather data on economy and industry| C
    C --> |Use tools like ratios and charts| D
    D --> |Interpret data for conclusions| E
    F --> |Periodic review and iteration| A
Explanation:
Step-1: Articulating Purpose and Context
Define the objectives of the analysis–such as assessing profitability, liquidity, or solvency. Identify stakeholders (e.g., investors, creditors, management) and tailor the analysis to their needs. Set the framework, including accounting standards (IFRS or US GAAP) and the time horizon (CFA Institute, 2025d).
Step-2: Collecting Input Data
Gather primary financial statements: income statement, balance sheet, and cash flow statement. Supplement this with industry benchmarks and macroeconomic data. Ensure the quality, accuracy, and completeness of all collected data (Investopedia, 2025a).
Step-3: Processing Data
Standardize data for comparability by adjusting for non-recurring items or differences in accounting policies. Compute financial ratios such as ROE, current ratio, and debt-to-equity. Use visualizations (e.g., charts, graphs) to uncover trends and patterns (Stickney et al., 2007).
Step-4: Analyzing and Interpreting Processed Data
Assess financial health by interpreting computed ratios. Benchmark against peer companies and industry averages. Identify strengths and weaknesses to determine strategic implications (Palepu et al., 2013).
Step-5: Developing and Communicating Conclusions
Summarize findings in a clear, concise report. Offer actionable recommendations–e.g., restructuring debt or improving efficiency. Tailor communication style and depth to fit the audience, whether board members, analysts, or external investors.
Step-6: Doing Follow-Up
Monitor outcomes of implemented actions and assess whether financial targets are met. Update the analysis regularly with new data and refine recommendations. Incorporate feedback to improve future analysis cycles.
Source:
CFA Program Curriculum’s Financial Reporting and Analysis readings covering ratio analysis, cash flow analysis, and IFRS/GAAP standards (CFA Institute, 2025d) alongside Investopedia’s overview of financial statement components (Investopedia, 2025a), Paul R. Brown’s strategic perspective on statement analysis and valuation (Stickney et al., 2007), and Palepu & Healy’s MBA-level treatment of business analysis and valuation using financial statements (Palepu et al., 2013).
Alternative Investments
***Alternative Investments:***
‘‘‘mermaid
graph TD
    A["Define Investment Objectives and Mandate"] --> B["Identify Alternative Asset Classes"]
    B --> C["Conduct Manager and Strategy Due Diligence"]
    C --> D["Perform Valuation and Pricing Analysis"]
    D --> E["Assess Risk and Liquidity"]
    E --> F["Allocate Alternatives in Portfolio"]
    F --> G["Monitor Performance and Rebalance"]
Explanation:
Step-1: Define Investment Objectives and Mandate – Clarify the purpose of including alternative investments–whether for diversification, higher return potential, or hedging against market volatility. Define constraints such as time horizon, liquidity needs, regulatory frameworks, and risk tolerance (CFA Institute, 2025a).
Step-2: Identify Alternative Asset Classes – Explore the universe of alternatives, including hedge funds, private equity, real estate, infrastructure, commodities, and venture capital. Assess how each class contributes to portfolio diversification via low correlation to traditional assets (Bodie et al., 2017; CAIA Association, 2025).
Step-3: Conduct Manager and Strategy Due Diligence – Evaluate managers based on their track record, investment philosophy, risk management, and operational quality. Understand the specific strategies (e.g., long/short, event-driven, global macro) and their alignment with investment mandates (CAIA Association, 2025; Metrick and Yasuda, 2010).
Step-4: Perform Valuation and Pricing Analysis – Address the unique valuation challenges of illiquid assets. Use models like discounted cash flow (DCF) or mark-to-model, and apply appropriate liquidity or opacity discounts. Compare performance with custom or market benchmarks (Metrick and Yasuda, 2010).
Step-5: Assess Risk and Liquidity – Identify key risks including market, manager, and operational risks. Analyze downside risk and tail event exposure. Evaluate liquidity risks, such as lock-up periods and redemption windows, that may affect rebalancing ability (CFA Institute, 2025a).
Step-6: Allocate Alternatives in Portfolio – Determine appropriate weighting of alternative assets, guided by expected return, volatility, and correlation with traditional investments. Make strategic allocation decisions with room for tactical adjustments based on market conditions (Bodie et al., 2017).
Step-7: Monitor Performance and Rebalance – Track returns over time, evaluate them against relevant benchmarks, and assess if performance remains consistent with expectations. Rebalance periodically to ensure alignment with objectives, risk profile, and current market landscape (CAIA Association, 2025).
Source:
CFA Program Curriculum’s Alternative Investments readings covering hedge funds, private equity, real assets, and due diligence frameworks (CFA Institute, 2025a)–together with Metrick & Yasuda’s deep dive into private equity and venture capital (Metrick and Yasuda, 2010), CAIA Association’s comprehensive CAIA-level materials on hedge funds, real estate, commodities, and other alternatives (CAIA Association, 2025), and Bodie, Kane & Marcus’s chapters on alternative asset classes and portfolio integration in Investments (Bodie et al., 2017).
Corporate Issuers
***Corporate Issuer Analysis:***
‘‘‘mermaid
graph TD
    A["Corporate Issuer Overview"] --> B["Industry Classification"]
    B --> C["Sector Trends and Competitive Landscape"]
    A --> D["Financial Statement Analysis"]
    D --> E["Profitability, Liquidity, Leverage"]
    A --> F["Credit Risk Assessment"]
    F --> G["Rating Agencies and Default Probabilities"]
    A --> H["Capital Structure and Issuance History"]
    H --> I["Bond Issuances and Debt Maturities"]
    A --> J["Corporate Governance and Management"]
    J --> K["Board Quality and Managerial Competence"]
    A --> L["Valuation and Investment Analysis"]
    L --> M["DCF, Relative Valuation, Multiples"]
Explanation:
Step-1: Corporate Issuer Overview – Begin with a high-level understanding of the firm’s business model, market positioning, and strategic objectives. This foundational context is essential for both equity and fixed income analysis (CFA Institute, 2025c).
Step-2: Industry Classification and Sector Trends – Classify the firm by sector or sub-sector (e.g., financials, consumer discretionary) and evaluate the competitive landscape. Analyze market trends, industry growth prospects, and systemic risks. This industry context shapes performance expectations and relative valuation (Penman, 2012).
Step-3: Financial Statement Analysis and Key Metrics – Analyze income statement, balance sheet, and cash flow data. Focus on metrics like revenue growth, operating margin, return on equity, and leverage. This step reveals the firm’s financial health and operational efficiency (Penman, 2012; CFA Institute, 2025c).
Step-4: Credit Risk Assessment and Rating Measures – Evaluate creditworthiness through agency ratings (e.g., S&P, Moody’s), credit spreads, and financial ratios. Analyze probability of default and credit cycle indicators. This step is vital for bondholders and fixed income portfolio managers (Fabozzi, 2012).
Step-5: Capital Structure, Issuance History, and Debt Profile – Examine the firm’s financing structure, including the mix of debt vs. equity, historical issuance patterns, and maturity schedules. This informs views on solvency and refinancing risks (Fabozzi, 2012).
Step-6: Corporate Governance and Leadership Quality – Assess governance practices such as board independence, shareholder rights, and disclosure quality. Evaluate the management team’s execution track record and alignment with shareholder interests (CFA Institute, 2025c).
Step-7: Valuation and Investment Analysis – Use valuation models like DCF, P/E, or EV/EBITDA to derive intrinsic value. Develop an investment thesis based on fundamental insights. These valuation techniques are central to both equity and credit investing (Penman, 2012).
Source:
CFA Program Curriculum’s Equity Investments and Fixed Income readings–which cover firm analysis, industry evaluation, and credit assessment frameworks (CFA Institute, 2025c)–along with Penman’s Financial Statement Analysis and Security Valuation for accounting-to-valuation linkages (Penman, 2012), and Fabozzi’s Bond Markets, Analysis, and Strategies for credit risk and corporate debt issuance insights (Fabozzi, 2012).
Appendix B Prompt Template
B.1 Structured Chain-of-Thought (ST-CoT)
ST-CoT for CFA Exam
You are a CFA (chartered financial analyst) taking a test to evaluate your knowledge of finance. You think step-by-step approach
to answer queries.
Follow these steps:
1. Think through the problem step by step within the <thinking> tags.
2. Provide your final, concise answer within the <output> tags.
The <thinking> sections are for your internal reasoning process only.
Do not include any part of the final answer in these sections.
The actual response to the query must be entirely contained within the <output> tags.
### Response Format:
<thinking>
[Reasoning through options A, B, and C to understand and solve the problem.]
</thinking>
<output>
"answer": [Final your answer (A , B , or C )]
</output>
B.2 FinCoT
FinCoT for CFA Exam
You are taking a test for the Chartered Financial Analyst ( CFA ) program designed to evaluate your knowledge of different topics in
finance. You think step-by-step approach with reflection to answer queries.
Follow these steps:
1. Think through the problem step by step reflect and verify while reasoning within the <thinking> tags.
2. Please and put the answer your final, concise answer within the <output> tags.
The <thinking> sections are for your internal reasoning process only.
Do not include any part of the final answer in these sections.
The actual response to the query must be entirely contained within the <output> tags.
Hint:{THOUGHT.get("embedding_expert_blueprints_[i]")}
### Response Format:
<thinking>
[Think step by step and respond with your thinking and the correct answer (A , B , or C ), considering the specific sector.]
</thinking>
<output>
"sector": [The sector being addressed],
"question": [The financial question],
"answer": [Reflect and verify the final answer (A , B , or C )]
</output>
B.3 Classify Domain
Classify Domain CFA Exam SYSTEM_INSTRUCTION = """You are a CFA expert. Categorize the given CFA question into exactly one of these categories:     Ethical and Professional Standards     - Code of Ethics, Standards of Professional Conduct, professional integrity     - Professional responsibilities, ethical decision-making, client interests     Category code: Ethics     Quantitative Methods     - Statistical analysis, probability theory, hypothesis testing     - Time value of money, financial mathematics, regression analysis     Category code: Quant.Meth.     Economic Analysis and Market Forces     - Microeconomics: supply, demand, market structures     - Macroeconomics: GDP, inflation, monetary policy, economic cycles     Category code: Economics     Financial Reporting and Analysis     - Financial statements, accounting standards, ratio analysis     - Balance sheets, income statements, cash flow analysis     Category code: Fin.Reporting     Corporate Finance and Issuers     - Capital structure, dividend policy, corporate governance     - Mergers & acquisitions, capital budgeting, risk management     Category code: Corp.Issuers     Equity Investments     - Stock valuation, equity markets, company analysis     - Market efficiency, equity portfolio management     Category code: EquityInvest.     Fixed Income Investments     - Bond markets, yield curves, duration analysis     - Credit analysis, fixed income portfolio management     Category code: FixedIncome     Derivative Instruments     - Options, futures, forwards, swaps     - Hedging strategies, derivative pricing, risk management     Category code: Derivatives     Alternative Investments     - Real estate, private equity, hedge funds     - Commodities, structured products, crypto assets     Category code: Alter.Invest.     Portfolio Management     - Asset allocation, portfolio construction, rebalancing     - Risk management, performance measurement, client objectives     Category code: Port.Manage. Respond with only the single most appropriate category code, nothing else. For example: Ethics, Port.Manage., etc. """
Appendix C Domain Distribution
Figure 6: GPT-4o classified the benchmark domain distribution of CFA. A random sample of 100 items was manually audited by a financial expert to validate domain labels.
Appendix D Average Output Tokens
Prompt Average Output Tokens (k)
Qwen2.5-7B
Qwen2.5-7B
Instruct
Qwen3-8B
Base
Qwen3-8B
Gemma-3-12B
IT
Qwen3-8B
Fin-R1
DianJin-R1
7B
Fin-o1-8B
SP 0.45 0.05* 0.89 0.32 0.27* 1.52 0.88 2.18 0.46*
UST-CoT 0.48 0.28 0.31* 0.46 0.39 1.50 0.58* 2.28 0.53
ST-CoT 0.39* 0.22 3.42 0.25* 0.31 1.35 2.22 7.20 0.58
FinCoT (All Blueprints) 2.22 0.29 0.38 0.36 0.32 1.23* 1.92 1.60* 0.79
Domain-wise performance of FinCoT
FinCoT (Economics) 0.36 0.38 0.99 0.39 0.38 1.25 2.01 12.65 0.76
FinCoT (FixedIncome) 0.42 0.27 4.55 0.30 0.32 1.24 2.31 8.31 0.81
FinCoT (Quant.Meth.) 0.48 0.27 3.07 0.31 0.35 1.22 2.17 8.60 0.80
FinCoT (EquityInvest.) 0.32 0.31 7.18 0.37 0.34 1.19 2.16 10.07 0.78
FinCoT (Port.Manage.) 0.38 0.26 0.56 0.30 0.33 1.20 2.14 9.46 0.79
FinCoT (Derivatives) 0.36 0.30 0.42 0.39 0.34 1.24 2.05 5.54 0.81
FinCoT (Fin. Reporting) 0.46 0.28 0.93 0.33 0.34 1.19 2.13 8.76 0.73
FinCoT (Alter.Invest.) 0.47 0.26 0.50 0.38 0.34 1.23 2.16 11.53 0.77
FinCoT (Corp.Issuers) 0.52 0.26 1.18 0.32 0.33 1.16 2.08 11.37 0.82
Table 2: Comparison of prompting techniques: average output token length (k) across models. Bold values highlight the prompt variant that uses the least tokens for each model. (*) Indicates that the change in average output token count among the model-level prompt variants is statistically significant ($p<0.05$) based on paired bootstrap testing; domain-specific rows are not tested for significance.
Appendix E Significance Testing
E.1 Accuracy
Model Baseline Comparison $\Delta$ (pp) 95% CI (pp) $p$-value Significant Qwen2.5-7B SP UST-CoT -13.76 [-15.89, -11.63] 0.0000 SP ST-CoT -16.27 [-18.60, -14.05] 0.0000 SP FinCoT (All) -7.94 [ -9.59, -6.30] 0.0000 UST-CoT ST-CoT -2.52 [ -3.49, -1.65] 0.0000 UST-CoT FinCoT (All) 5.81 [ 4.46, 7.27] 0.0000 ST-CoT FinCoT (All) 8.33 [ 6.69, 9.98] 0.0000 Qwen2.5-7B-Instruct SP UST-CoT 6.00 [ 4.65, 7.56] 0.0000 SP ST-CoT 4.84 [ 3.59, 6.20] 0.0000 SP FinCoT (All) 4.55 [ 3.29, 5.91] 0.0000 UST-CoT ST-CoT -1.16 [-1.84, -0.58] 0.0000 UST-CoT FinCoT (All) -1.45 [-2.23, -0.78] 0.0000 ST-CoT FinCoT (All) -0.29 [-0.68, 0.00] 0.1024 Qwen3-8B-Base SP UST-CoT -9.40 [-11.24, -7.66] 0.0000 SP ST-CoT -15.31 [-17.54,-13.18] 0.0000 SP FinCoT (All) -17.35 [-19.77,-15.02] 0.0000 UST-CoT ST-CoT -5.91 [-7.36, -4.55] 0.0000 UST-CoT FinCoT (All) -7.96 [-9.59, -6.30] 0.0000 ST-CoT FinCoT (All) -2.04 [-3.00, -1.26] 0.0000 Qwen3-8B SP UST-CoT 7.95 [ 6.30, 9.69] 0.0000 SP ST-CoT 6.60 [ 5.14, 8.14] 0.0000 SP FinCoT (All) 6.70 [ 5.23, 8.24] 0.0000 UST-CoT ST-CoT -1.35 [-2.13, -0.68] 0.0000 UST-CoT FinCoT (All) -1.26 [-1.94, -0.68] 0.0000 ST-CoT FinCoT (All) 0.10 [ 0.00, 0.29] 0.7370 Qwen3-8B (Thinker) SP UST-CoT -0.87 [-1.45, -0.39] 0.0002 SP ST-CoT 0.00 [ 0.00, 0.00] 2.0000 SP FinCoT (All) 0.96 [ 0.39, 1.55] 0.0002 UST-CoT ST-CoT 0.87 [ 0.39, 1.45] 0.0002 UST-CoT FinCoT (All) 1.83 [ 1.07, 2.71] 0.0000 ST-CoT FinCoT (All) 0.96 [ 0.39, 1.55] 0.0002 Gemma-3-12B-IT SP UST-CoT -24.999 [-27.71, -22.38] 0.0000 SP ST-CoT -23.934 [-26.55, -21.32] 0.0000 SP FinCoT (All) -22.765 [-25.39, -20.16] 0.0000 UST-CoT ST-CoT 1.065 [ 0.48, 1.74] 0.0002 UST-CoT FinCoT (All) 2.234 [ 1.36, 3.20] 0.0000 ST-CoT FinCoT (All) 1.169 [ 0.58, 1.84] 0.0000 Fin-R1 SP UST-CoT -9.49 [-11.34, -7.75] 0.0000 SP ST-CoT -8.62 [-10.37, -6.88] 0.0000 SP FinCoT (All) -10.07 [-11.92,-8.24] 0.0000 UST-CoT ST-CoT 0.87 [ 0.39, 1.45] 0.0002 UST-CoT FinCoT (All) -0.58 [-1.07, -0.19] 0.0042 ST-CoT FinCoT (All) -1.45 [-2.23, -0.78] 0.0000 Dianjin-R1-7B SP UST-CoT 10.662 [ 8.82, 12.60] 0.0000 SP ST-CoT 9.594 [ 7.75, 11.43] 0.0000 SP FinCoT (All) -1.361 [ -2.13, -0.68] 0.0000 UST-CoT ST-CoT -1.068 [ -1.74, -0.48] 0.0000 UST-CoT FinCoT (All) -12.023 [-14.05, -10.08] 0.0000 ST-CoT FinCoT (All) -10.956 [-12.89, -9.01] 0.0000 Fino1-8B SP UST-CoT 0.294 [ 0.00, 0.68] 0.0984 SP ST-CoT 1.264 [ 0.58, 1.94] 0.0000 SP FinCoT (All) 2.429 [ 1.55, 3.39] 0.0000 UST-CoT ST-CoT 0.970 [ 0.39, 1.55] 0.0000 UST-CoT FinCoT (All) 2.134 [ 1.26, 3.00] 0.0000 ST-CoT FinCoT (All) 1.165 [ 0.58, 1.84] 0.0000
Table 3: Paired bootstrap significance testing (B=10, [^0] B=10{,}000 samples) for accuracy differences across prompt strategies. $\Delta$ indicates average accuracy difference (in percentage points), with 95% confidence intervals (CI) and $p$-values. A result is considered statistically significant if $p<0.05$.
E.2 Average Output Tokens
Model Baseline Comparison $\Delta$ (k) 95% CI (k) $p$-value Significant Qwen2.5-7B SP UST-CoT -0.09867 [-0.09867, -0.09867] 0.0000 SP ST-CoT 0.02273 [ 0.02273, 0.02273] 0.0000 SP FinCoT (All) -0.11555 [-0.11555, -0.11555] 0.0000 UST-CoT ST-CoT 0.12140 [ 0.12140, 0.12140] 0.0000 UST-CoT FinCoT (All) -0.01688 [-0.01688, -0.01688] 0.0000 ST-CoT FinCoT (All) -0.13828 [-0.13828, -0.13828] 0.0000 Qwen2.5-7B-Instruct SP UST-CoT -0.227 [-0.227, -0.227] 0.0000 SP ST-CoT -0.169 [-0.169, -0.169] 0.0000 SP FinCoT (All) -0.241 [-0.241, -0.241] 0.0000 UST-CoT ST-CoT 0.058 [ 0.058, 0.058] 0.0000 UST-CoT FinCoT (All) -0.014 [-0.014, -0.014] 0.0000 ST-CoT FinCoT (All) -0.072 [-0.072, -0.072] 0.0000 Qwen3-8B-Base SP UST-CoT -0.584 [-0.584, -0.584] 0.0000 SP ST-CoT 2.522 [ 2.522, 2.522] 0.0000 SP FinCoT (All) -0.516 [-0.516, -0.516] 0.0000 UST-CoT ST-CoT 3.106 [ 3.106, 3.106] 0.0000 UST-CoT FinCoT (All) 0.068 [ 0.068, 0.068] 0.0000 ST-CoT FinCoT (All) -3.038 [-3.038, -3.038] 0.0000 Qwen3-8B SP UST-CoT 0.141 [ 0.141, 0.141] 0.0000 SP ST-CoT -0.067 [-0.067, -0.067] 0.0000 SP FinCoT (All) 0.048 [ 0.048, 0.048] 0.0000 UST-CoT ST-CoT -0.208 [-0.208, -0.208] 0.0000 UST-CoT FinCoT (All) -0.093 [-0.093, -0.093] 0.0000 ST-CoT FinCoT (All) 0.115 [ 0.115, 0.115] 0.0000 Qwen3-8B (Thinker) SP UST-CoT -0.018 [-0.018, -0.018] 0.0000 SP ST-CoT -1.271 [-1.271, -1.271] 0.0000 SP FinCoT (All) -0.168 [-0.168, -0.168] 0.0000 UST-CoT ST-CoT -1.253 [-1.253, -1.253] 0.0000 UST-CoT FinCoT (All) -0.150 [-0.150, -0.150] 0.0000 ST-CoT FinCoT (All) 1.103 [ 1.103, 1.103] 0.0000 Gemma-3-12B-IT SP UST-CoT 0.11985 [ 0.11985, 0.11985] 0.0000 SP ST-CoT 0.03661 [ 0.03661, 0.03661] 0.0000 SP FinCoT (All) 0.04523 [ 0.04523, 0.04523] 0.0000 UST-CoT ST-CoT -0.08324 [-0.08324, -0.08324] 0.0000 UST-CoT FinCoT (All) -0.07462 [-0.07462, -0.07462] 0.0000 ST-CoT FinCoT (All) 0.00862 [ 0.00862, 0.00862] 0.0000 Fin-R1 SP UST-CoT 1.526 [ 1.526, 1.526] 0.0000 SP ST-CoT 1.338 [ 1.338, 1.338] 0.0000 SP FinCoT (All) 1.035 [ 1.035, 1.035] 0.0000 UST-CoT ST-CoT -0.188 [-0.188, -0.188] 0.0000 UST-CoT FinCoT (All) -0.491 [-0.491, -0.491] 0.0000 ST-CoT FinCoT (All) -0.303 [-0.303, -0.303] 0.0000 Dianjin-R1-7B SP UST-CoT 0.10023 [ 0.10023, 0.10023] 0.0000 SP ST-CoT 5.02159 [ 5.02159, 5.02159] 0.0000 SP FinCoT (All) -0.57669 [-0.57669, -0.57669] 0.0000 UST-CoT ST-CoT 4.92136 [ 4.92136, 4.92136] 0.0000 UST-CoT FinCoT (All) -0.67692 [-0.67692, -0.67692] 0.0000 ST-CoT FinCoT (All) -5.59828 [-5.59828, -5.59828] 0.0000 Fino1-8B SP UST-CoT 0.06788 [ 0.06788, 0.06788] 0.0000 SP ST-CoT 0.11765 [ 0.11765, 0.11765] 0.0000 SP FinCoT (All) 0.33273 [ 0.33273, 0.33273] 0.0000 UST-CoT ST-CoT 0.04977 [ 0.04977, 0.04977] 0.0000 UST-CoT FinCoT (All) 0.26485 [ 0.26485, 0.26485] 0.0000 ST-CoT FinCoT (All) 0.21508 [ 0.21508, 0.21508] 0.0000
Table 4: Paired bootstrap significance testing (B=10, [^0] B=10{,}000 samples) for average output token differences across prompt strategies. $\Delta$ indicates mean difference in output length (in thousands of tokens), with 95% confidence intervals and $p$-values. A result is significant if $p<0.05$.
Appendix F Radar Behavior Accuracy
Figure 7: Overall FinCoT behaviour accuracy.
(a) Qwen2.5-7B
(b) Qwen2.5-7B-Instruct
(c) Qwen3-8B-Base
(d) Qwen3-8B
Figure 8: Radar charts for each model variant.
(a) Qwen3-8B (Thinker)
(b) Gemma-3-12B-IT
(c) Fin-R1
(d) Dianjin-R1-7B
Figure 9: Radar charts for each model variant (charts 5–8).
Figure 10: Radar charts for each model variant (charts 9).